OpenStudy (samsan9):

How to solve tan 195 using the sum and difference formula?

terenzreignz (terenzreignz):

With magic. D'uh :D Or... you could find two angles whose tangents you DO know, (say, 30 or 45) and that add up to 195. Can you find two such angles?

OpenStudy (samsan9):

its is 135 and 60 i believe?

terenzreignz (terenzreignz):

Great ^_^ What are tan(135) and tan(60) ?

OpenStudy (samsan9):

i got 1 for tan 135 and (square root)3/3 for 60

terenzreignz (terenzreignz):

You might want to take a better look at tan(135) ... ^_^

OpenStudy (samsan9):

is it -1

terenzreignz (terenzreignz):

Yes... lol it's in the second quadrant, where ONLY sine and cosecant are positive, the rest should be negative. What about tan(60) though?

OpenStudy (samsan9):

i apoligize is it square root of 3?

terenzreignz (terenzreignz):

Yes, it is ^_^ \[\large \tan(135^o) = -1\\\large \tan(60^o) = \sqrt 3\] Now, we have.. \[\Large \tan(195^o) = \tan(\color{blue}{135^o} + \color{red}{60^o})\] And the formula: \[\Large \tan(\color{blue}\alpha + \color{red}\beta ) = \frac{\tan(\color{blue}\alpha)+\tan(\color{red}\beta)}{1-\tan(\color{blue}\alpha)\tan(\color{red}\beta)}\] Just plug in... and be done ;)

OpenStudy (samsan9):

is it -1+(squareroot)3

OpenStudy (samsan9):

over 1+\[1+\sqrt{3}\]

terenzreignz (terenzreignz):

This? \[\Large \frac{-1+\sqrt 3}{1+\sqrt3}\]

OpenStudy (samsan9):

yes

terenzreignz (terenzreignz):

It's correct. Well done ^_^

OpenStudy (samsan9):

is it -4+2(squareroot)3/-2?

terenzreignz (terenzreignz):

You rationalise denominators? lol We usually leave that as is... Anyway, you're right, but you might as well multiply both the numerator and denominator by -1 so as to remove that ugly negative sign down below: \[\Large \frac{4-2\sqrt3}{2}\] And then just simplify ^_^

OpenStudy (samsan9):

so is it 2-squareroot(3)

terenzreignz (terenzreignz):

That is correct :)